Sandras father is 45 years old this year. His age is four times more than Sandras age. How old is Sandra this year?
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

Sandras father is 45 years old this year. His age is four times more than Sandras age. How old is Sandra this year?
--------
She's 9
45/5 = 9

Answer by greenestamps(13200)



"...four times more than Sandra's age..." means Sandra's age, plus her age four more times -- i.e., it means five times AS OLD AS Sandra.

If x is Sandra's age, then
5x = 45
x = 45/5 = 9

Sandra's age is 9.